# Quantum Otto engine mimicking Carnot near pseudotransitions in the 1D extended Hubbard model in the atomic limit

### v1

@@ I Introduction
In the past two decades, significant efforts have been made to explore thermodynamic processes where quantum features of matter are crucial. This research has successfully combined theoretical proposals [1] with experimental evidence [2], bridging the gap between quantum thermodynamics concepts and practical applications. Quantum heat engines, like Carnot and Otto, have been studied, with their characteristics compared to classical counterparts. Additionally, a quantum heat engine model has been analyzed for efficiency and irreversibility [3, 4]. Quantum thermodynamics offers a framework for understanding and optimizing engines and refrigerators [5]. The study of quantum heat engines based on harmonic oscillators has enhanced our understanding of optimization and irreversibility [6, 7].
@@
Several quantum machines have been explored using one-dimensional models. In this context, a quantum Otto heat engine based on a multiferroic chain as the working substance was investigated in [18]. Halpern et al. [19] proposed harnessing many-body localization for thermodynamic tasks, designing an Otto engine with mesoscale parallel reliability. In [20], the thermodynamic cycle of an Otto engine with a quantum Ising chain as the working substance was examined. This cycle involves sweeps of the transverse magnetic field in thermal isolation, alternated with thermalization strokes involving reservoirs at different temperatures. The performance of quantum thermal machines using many-body quantum systems or spin-chain models, periodically connected to external baths via local couplings, was discussed in [21]. More recently, the spin-chain-star shape model was analyzed as the working substance of a quantum Otto cycle in [22]. Additionally, researchers have studied a finite-time quantum Otto cycle in a 1D Bose gas, analyzing efficiency-power trade-offs [23]. Nautiyal et al. [24] examined an imperfect quantum Otto engine, evaluating work, power, and optimal operating points under less ideal conditions.
@@ II.3 Frustrated and alternating pair phase transition
Here we focus in two ground state phases at half-filling band $\rho=1$. The first phase is the frustrated phase $FR_{2}$ given by
@@
The AP phase has no residual entropy and can be identified also as a charge density wave (CDW)[29].
@@ II.5 Magnetic Properties
In Fig. 1, the isentropic curve is illustrated in the $h-T$ plane, considering the following fixed parameters: $U=10$, $V=5.4$, $\mu=15.8$. The isentropic curve exhibits unusual behavior over a certain range of entropy, as shown in the legend. Specifically, for $0.40.3\apprle h<$ 0.4, all isentropic curves are nearly overlapping. This unusual behavior significantly affects the thermal properties and, for lower magnetic fields, the curves remain densely packed, highlighting the influence of the anomalous region. The blue curves delineate the quasi-alternating pair ($qAP$) region for $h\lesssim 0.4$, whereas for $h\gtrsim 0.4$, the area below the blue curve corresponds to the quasi-paramagnetic ($qPM$) region. Additionally, for a null magnetic field and at temperatures roughly around $T\sim 0.5-0.6$, there is evidence of a strong change in entropy, indicating a transition from the $qAP$ to the $qFR_{2}$ region. In Fig. 1, below blue curve ($T\lesssim 0.01$) basically the entropy is null, see Fig. 2a from another perspective.
@@
Figure 2: (a) Entropy as a function of temperature $T$, with parameters $U=10$, $V=5.3$ and $\mu=15.8$. (b) Entropy variation $\Delta\mathcal{S}=\mathcal{S}(T,h_{1})-\mathcal{S}(T,0)$ as a function of temperature $T$, with fixed parameters $U=10$, $\mu=15.8$, and $V=5.3$, for $h_{1}=0.29$ (violet curve). Similarly, for $h_{2}=0.38$ (green curve) and fixed parameters $U=10$, $\mu=15.8$, and $V=5.4$. (c) Adiabatic temperature change $\Delta T_{ad}$ as a function of temperature, for the same set of parameters as in (b). (d) The product $\Delta\mathcal{S}\Delta T_{ad}$ as a function of temperature, under the same conditions as in (b).
@@ III.2 Operation modes of machine
Considering the one-dimensional EHM in the atomic limit as the working substance, the work is linked to changes in energy levels as the magnetic field $h$ is altered. By adjusting the parameters of the reversible cycle in a 1D EHM in the atomic limit, the total work of the cycle can be reversed in accordance with the second law of thermodynamics, which, as formulated by Clausius, defines operational regions for thermodynamically reversible cycles based on temperature differences ($T_{h}>T_{l}$). In heat engine mode, the system absorbs heat from a hotter bath at $T_{h}$ ($Q_{in}>0$) and releases it to a cooler bath at $T_{l}$ ($Q_{out}<0$), partially converting the heat into work ($W_{net}>0$), as shown in Table 1. In refrigerator mode, the model induces heat flow from the cooler bath at $T_{l}$ to the hotter bath at $T_{h}$, requiring more work input than work output, with net work performed ($W_{net}<0$). As a thermal accelerator, the system uses net work ($W_{net}<0$) to increase heat flow from the hotter to the cooler bath. In heater mode, the net work ($W_{net}<0$) is applied to generate heat flow in both baths, causing the 1D EHM in the atomic limit to release heat in both baths ($Q_{in}<0$ and $Q_{out}<0$). All these processes are summarized in Table 1. In particular, we apply the Carnot and Otto engine processes as detailed in Appendix A.
@@ IV Results of Carnot and Otto machine
Figure 4: Heat absorbed ($Q_{in}$), Released heat ($Q_{out}$), and Work performed ($W_{net}$) as a function of Magnetic field, assuming $U=10$, $V=5.4$, $\mu=15.8$. Particularly here we fixed the $T_{h}=0.56$ and $T_{l}=0.28$.
@@
Figure 5: Thermal efficiency of Carnot and Otto engine as a function of magnetic field, considering $U=10$, $V=5.4$, $\mu=15.8$. Particularly here we fixed the $T_{h}=0.56$ and $T_{l}=0.28$.
@@
Figure 7: Thermal efficiency of Carnot and Otto engine as a function of magnetic field, considering $U=10$, $V=5.3$, $\mu=15.8$. Particularly here we fixed the $T_{h}=0.43$ and $T_{l}=0.19$.
@@ V Conclusion
Figs. 1 and 2 illustrate that the pseudo-transition region offers a unique thermodynamic regime with smooth entropy variation, enabling efficient quantum thermodynamic cycles. Unlike true critical points, which require fine-tuning or result in divergences, the pseudo-transition region provides stable thermodynamic control. These findings suggest that engineered quantum materials exhibiting pseudo-transitions could be promising candidates for quantum heat engines and other thermodynamic applications.

### v2

# Quantum Otto engine mimicking Carnot near pseudotransitions in the one-dimensional extended Hubbard model in the atomic limit

@@ I Introduction
In the past two decades, significant efforts have been made to explore thermodynamic processes where quantum features of matter are crucial. This research has successfully combined theoretical proposals [1] with experimental evidence [2], bridging the gap between quantum thermodynamics concepts and practical applications. Quantum heat engines, like Carnot and Otto, have been studied, and their characteristics have been compared to classical counterparts. Additionally, a quantum heat engine model has been analyzed for efficiency and irreversibility [3, 4]. Quantum thermodynamics offers a framework for understanding and optimizing engines and refrigerators [5]. The study of quantum heat engines based on harmonic oscillators has enhanced our understanding of optimization and irreversibility [6, 7].
@@
Several quantum machines have been explored using one-dimensional models. In this context, a quantum Otto heat engine based on a multiferroic chain as the working substance was investigated in [18]. Halpern et al. [19] proposed harnessing many-body localization for thermodynamic tasks, designing an Otto engine with mesoscale parallel reliability. In [20], the thermodynamic cycle of an Otto engine with a quantum Ising chain as the working substance was examined. This cycle involves sweeps of the transverse magnetic field in thermal isolation, alternated with thermalization strokes involving reservoirs at different temperatures. The performance of quantum thermal machines using many-body quantum systems or spin-chain models, periodically connected to external baths via local couplings, was discussed in [21]. More recently, the spin-chain-star shape model has been analyzed as the working substance of a quantum Otto cycle in [22]. Additionally, researchers have studied a finite-time quantum Otto cycle in a 1D Bose gas, analyzing efficiency-power trade-offs [23]. Nautiyal et al. [24] examined an imperfect quantum Otto engine, evaluating work, power, and optimal operating points under less ideal conditions.
@@ II.3 Frustrated and alternating pair phase transition
Here we focus on two ground state phases at half-filling band $\rho=1$. The first phase is the frustrated phase $FR_{2}$ given by
@@
The AP phase has no residual entropy and can be identified as a charge density wave (CDW)[29].
@@ II.5 Magnetic Properties
In Fig. 1, the isentropic curve is illustrated in the $h-T$ plane, considering the following fixed parameters: $U=10$, $V=5.4$, and $\mu=15.8$. The isentropic curve exhibits unusual behavior over a certain range of entropy, as shown in the legend. Specifically, for $0.40.3\apprle h<$ 0.4, all isentropic curves are nearly overlapping. This unusual behavior significantly affects the thermal properties and, for lower magnetic fields, the curves remain densely packed, highlighting the influence of the anomalous region. The blue curves delineate the quasi-alternating pair ($qAP$) region for $h\lesssim 0.4$, whereas for $h\gtrsim 0.4$, the area below the blue curve corresponds to the quasi-paramagnetic ($qPM$) region. Additionally, for a null magnetic field and at temperatures roughly around $T\sim 0.5-0.6$, there is evidence of a strong change in entropy, indicating a transition from the $qAP$ to the $qFR_{2}$ region. In Fig. 1, below blue curve ($T\lesssim 0.01$) basically the entropy is null, see Fig. 2a from another perspective.
@@
Figure 2: (a) Entropy as a function of temperature $T$, with parameters $U=10$, $V=5.3$ and $\mu=15.8$. (b) Entropy variation $\Delta\mathcal{S}=\mathcal{S}(T,h_{1})-\mathcal{S}(T,0)$ as a function of temperature $T$, with fixed parameters $U=10$, $\mu=15.8$, and $V=5.3$, for $h_{1}=0.29$ (violet curve). Similarly, for $h_{1}=0.38$ (green curve) and fixed parameters $U=10$, $\mu=15.8$, and $V=5.4$. (c) Adiabatic temperature change $\Delta T_{ad}$ as a function of temperature, for the same set of parameters as in (b). (d) The product $\Delta\mathcal{S}\Delta T_{ad}$ as a function of temperature, under the same conditions as in (b).
@@ III.2 Operation modes of machine
Considering the one-dimensional EHM in the atomic limit as the working substance, the work is linked to changes in energy levels as the magnetic field $h$ is altered. By adjusting the parameters of the reversible cycle in a 1D EHM in the atomic limit, the total work of the cycle can be reversed in accordance with the second law of thermodynamics, which, as formulated by Clausius, defines operational regions for thermodynamically reversible cycles based on temperature differences ($T_{h}>T_{l}$). In heat engine mode, the system absorbs heat from a hotter bath at $T_{h}$ ($Q_{in}>0$) and releases it to a cooler bath at $T_{l}$ ($Q_{out}<0$), partially converting the heat into work ($W_{net}>0$), as shown in Table 1. In refrigerator mode, the model induces heat flow from the cooler bath at $T_{l}$ to the hotter bath at $T_{h}$, requiring more work input than work output, resulting in net work performed ($W_{net}<0$). As a thermal accelerator, the system uses net work ($W_{net}<0$) to increase heat flow from the hotter to the cooler bath. In heater mode, the net work ($W_{net}<0$) is applied to generate heat flow in both baths, causing the 1D EHM in the atomic limit to release heat in both baths ($Q_{in}<0$ and $Q_{out}<0$). All these processes are summarized in Table 1. In particular, we apply the Carnot and Otto engine processes as detailed in Appendix A.
@@ IV Results of Carnot and Otto machine
Figure 4: Heat absorbed ($Q_{in}$), released heat ($Q_{out}$), and work performed ($W_{net}$) as a function of magnetic field, assuming $U=10$, $V=5.4$, and $\mu=15.8$. Particularly here we fixed the $T_{h}=0.56$ and $T_{l}=0.28$.
@@
Figure 5: Thermal efficiency of Carnot and Otto engine as a function of magnetic field, considering $U=10$, $V=5.4$, and $\mu=15.8$. Particularly here we fixed the $T_{h}=0.56$ and $T_{l}=0.28$.
@@
Figure 7: Thermal efficiency of Carnot and Otto engine as a function of magnetic field, considering $U=10$, $V=5.3$, and $\mu=15.8$. Particularly here we fixed the $T_{h}=0.43$ and $T_{l}=0.19$.
@@ V Conclusion
Figs. 1 and 2 illustrate that the pseudo-transition region offers a unique thermodynamic regime characterized smooth entropy variation, enabling efficient quantum thermodynamic cycles. Unlike true critical points, which require fine-tuning or result in divergences, the pseudo-transition region provides stable thermodynamic control. These findings suggest that engineered quantum materials exhibiting pseudo-transitions could be promising candidates for quantum heat engines and other thermodynamic applications.